# Robust Signal Detection with Quadratically Convex Orthosymmetric Constraints

### v2

@@ 3.1 Lower Bounds
In this section, we introduce three lower bounds, i.e., Theorem 3.2, 3.3, and 3.4, which independently work on determining the lower bound of $\rho_{\text{critical}}$.
@@ 3.1.1 Based on a Geometric Property
The following lemma is an important basis for the theoretical lower bounds in Theorems 3.2 and 3.3. Exploiting the properties of QCO sets and the definition of the Kolmogorov widths, Lemma 3.1 outputs a vector with large $\ell_{2}$ norm but small $\ell_{\infty}$ norm, which is fundamental for the prior of $v$ constructed under $H_{1}$. The proof, which involves an application of the minimax theorem, is provided in Appendix B.1.
@@
We now are ready to state the first main lower bound.
Theorem 3.2.
@@
Moreover, one possible explicit expression of $c(\alpha)$ is
@@
Theorem 3.2 above states that for fixed $\alpha$, if $\rho\lesssim\frac{(k_{1}^{\star})^{1/4}}{\sqrt{N}}\sigma$, it is impossible to find a test that performs well under $H_{0}$ and all possible $H_{1}$ uniformly. As a corollary, this result implies that it is also impossible to distinguish between $H_{0}$ and $H_{1}$ in the presence of an adversary $\mathcal{C}$ under the above condition, since a simple strategy for $\mathcal{C}$ is to keep all samples untouched. The proof of Theorem 3.2 relies on the selected vector in Lemma 3.1 and Le Cam’s method of “fuzzy hypothesis”.
Moreover, it is worth noting the marginal case when $k_{1}^{\star}=0$. The definition of $k_{1}^{\star}$ only relies on $K$, NN, $\sigma$, and does not depend on the original samples $\widetilde{\mathbf{X}}$ or the observations $\mathbf{X}$. Therefore, we shall assume in the following text that $D_{0}(K)=\sup\limits_{\theta\in K}\left\lVert\theta\right\rVert_{2}>\frac{1}{\sqrt{N}}\sigma$, and therefore $k_{1}^{\star}\geq 1$. Otherwise, the test will not be feasible in the first place. We kindly refer the readers to Appendix B.2 for the proof of this argument and Theorem 3.2.
Recall that for uncontaminated observations without constraint, the critical value of the $\ell_{2}$ norm is $\rho_{\text{critical}}\asymp\frac{d^{1/4}}{\sqrt{N}}\sigma$ (see Section 3.3.6 in Ingster and Suslina, (2003) for more details). Theorem 3.2 generalizes this classic result in the sense that additional prior knowledge the mean is allowed and the lower bound is refined as a consequence. Indeed, Theorem 3.2 recovers the classic result when we set $\mathcal{X}=\mathbb{R}^{d}$ (where $k_{1}^{\star}=d$ according to Example 2.5). Moreover, the theorem implies that when we have non-trivial constraint $K$ on $\mu$, the essential dimensional attribute of the problem is the finite $k_{1}^{\star}$ instead of $d$. (Recall that $k_{1}^{\star}<\infty$ is ensured even if $K\subset\ell_{2}$.)
@@ 3.1.2 Based on the Framework of Differential Privacy
The concept of differential privacy was first introduced by Dwork et al., (2006) in the study of private data analysis. It was later introduced to the goodness-of-fit test (also called identity test) in statistics. See Gaboardi et al., (2016), Rogers and Kifer, (2017), Cai et al., (2017) for some earlier works of this application. Differential privacy requires a random algorithm to be robust against the disturbance of the input. Please see Section B.3 for the formal definition of differential privacy. In Theorem 3.3, we will combine Lemma 3.1 with an argument of optimal coupling and results from differential privacy to obtain another requirement for $\rho$. Please refer to Section B.3 for the techniques and proofs.
Theorem 3.3.
For the testing problem 1.2 with an adversary $\mathcal{C}$ from strong $\epsilon$-contamination model 2.8, it is impossible to derive a test that achieves type-I error and type-II error both less than 0.1 if $\rho\lesssim\frac{(k_{2}^{\star})^{1/4}\sqrt{\epsilon}\sigma}{N^{1/4}}$. Formally, if $\rho\lesssim\frac{(k_{2}^{\star})^{1/4}\sqrt{\epsilon}\sigma}{N^{1/4}}$, then we have
@@
where $A_{s}(0.1)$ is the set of acceptable tests under $0.1$ type-I error defined as (1.3).
We use $0.1$ for convenience, but it can be replaced by any unversal sufficiently small constant. It is worth noting that in the high-dimensional setting, where $d$, and possibly $k_{2}^{\star}$ can be substantially larger than $N$, one should be cautious that $\frac{(k_{2}^{\star})^{1/4}\sqrt{\epsilon}\sigma}{N^{1/4}}$ shall not exceeds $D_{0}(K)$, which is the maximal possible $\ell_{2}$ norm of the vector in $K$. The desired test will be impossible if this is not achieved.
@@ 3.1.3 Based on the Estimation Lower Bounds
The final lower bound, as established in Theorem 3.4 below, contrasts with Theorems 3.2 and 3.3 in that it is determined solely on the diameter of $K$ (which is $D_{0}(k)$), the fraction of contamination $\epsilon$, and the scale of the noise $\sigma$ while remaining independent of $N$.
Theorem 3.4.
For the signal detection problem (1.2), if $\epsilon\gtrsim\frac{1}{\sqrt{N}}$ and $\rho\lesssim\epsilon\sigma$, we always have
$\inf\limits_{\phi\in A_{s}}\sup\limits_{\mu\in K,\left\lVert\mu\right\rVert_{2}\geq\rho}\sup\limits_{\mathcal{C}}\mathbb{P}_{\mu}(\phi(\mathcal{C}(\widetilde{\mathbf{X}}))=0)\gtrsim 1,$ (3.1)
@@
The condition $\epsilon\gtrsim\frac{1}{\sqrt{N}}$ can be safely removed. See our discussions in Section 3.3.
Recall that $D_{0}(K)=\sup\limits_{\theta\in K}\left\lVert\theta\right\rVert_{2}$ is the supremum of the distance between a point in $K$ and the origin. (3.1) implies that to find a test with type-I and type-II error uniformly (over the adversary $\mathcal{C}$ and the mean $\mu$) less than $\alpha$, we should have either $\rho\gtrsim\epsilon\sigma$ or $\epsilon=\mathcal{O}(1/\sqrt{N})$. Since we always have $\left\lVert\mu\right\rVert_{2}\leq D_{0}(K)$, consequently, if $\epsilon\gtrsim 1/\sqrt{N}$ and $\epsilon\sigma\gtrsim D_{0}(K)$, the desired test will be theoretically impossible.
The proof of Theorem 3.4 is motivated by a similar lower bound established for the corresponding estimation problem under star-shaped constraints by Prasadan and Neykov, (2025). It is based on a careful construction of a mixture distribution which is represented in two different ways, and a computation of the total variation between them. The full proof is provided in Section B.4. This bound aligns with intuition, as testing is generally easier than estimation. See Section 5 for further details.
@@ 3.2.1 Theoretical Algorithm
An intuitive strategy for handling the adversary is to identify subsets of $\mathbf{X}$ that exhibit a consistency property. As we will see later, such subsets behave as if they consisted entirely of authentic samples with high probability.
@@
The following theorem establishes the existence of a valid testing procedure under condition (3.2). Specifically, we demonstrate that, when (3.2) holds, there exists a desired test $\phi_{e}$ and at least one consistent subset $S$ with respect to $\phi_{e}$, upon which the entire testing task becomes feasible. To achieve this, we explicitly define an event $E_{e}$ of $\chi^{2}$-type, on which $\phi_{e}$ is determined. Furthermore, we prove that the set of original samples $\widetilde{\mathbf{X}}$, though not observed, is consistent with high probability. Finally, we show that the proposed test $\phi_{e}$ achieves uniformly small type-I and type-II errors. The detailed proof is provided in Section B.5.
Theorem 3.6.
Given a dataset $\mathbf{X}$ from the strong $\epsilon$-contamination model (see Definition 2.8) and a fixed error tolerance $\alpha$, for any $k\in[1,d]$ [^2] When $K\subset\ell_{2}$, we allow $1\leq k<\infty$., there exists a test $\phi_{e}$ for (1.2) such that the type-I and type-II errors are both uniformly less than $\alpha$ if $\rho^{2}\gtrsim E_{\text{raw}}^{2}(\epsilon,K,N,\sigma^{2},k)$ where $E_{\text{raw}}^{2}(\epsilon,K,N,\sigma^{2},k)$ is defined as
@@
We can further leverage the results of Theorem 3.6 by the definition of $k_{1}^{\star}$ and $k_{2}^{\star}$ and optimizing over $k$.
@@
Corollary 3.7 actually selects the smaller optimal dimension between $k_{1}^{\star}$ and $k_{2}^{\star}$ and consequently the corresponding optimal projection, based on which $\phi_{e}$ is performed. We will show in Section 3.3 that the upper bound (3.3) does nearly match the theoretical lower bounds in the sense that it only has a redundant $\ln\left(\frac{1}{\epsilon}\right)$ factor.
A significant improvement implied by Theorem 3.6 is that it does not require $\epsilon\rightarrow 0$, which is an essential assumption in many previous works. However, the test $\phi_{e}$ is limited in that its construction requires a full scan of all subsets of $\mathbf{X}$, which consumes exponential time with respect to $N$. This prohibitive computational cost prevents it from being practical when $N$ is large. In the next section, motivated by the filtering method given by Canonne et al., (2023), we propose an algorithm with polynomial-degree time complexity when the prior constraint $K$ and adversary $\mathcal{C}$ present.
3.2.2 Polynomial Time Algorithm
We will develop a polynomial time algorithm [^3] Here when we say “polynomial time algorithm”, we do assume that calculating the optimal projections (resulting in the optimal dimensions $k_{1}^{\star}$ and $k_{2}^{\star}$) can be done in polynomial time — as is the case for any ellipsoid or hyperrectangle. in this section to detect whether $\mu=0$ or not while requiring a slightly stronger condition compared to the bound (3.2). Specifically, we will prove the following theorem.
Theorem 3.8.

### v3

@@ 3.1 Lower Bounds
In this section, we introduce three independent lower bounds, i.e., Theorem 3.2, 3.5, and 3.6, which jointly and fully determine the lower bound of $\rho_{\text{critical}}$.
@@ 3.1.1 Based on a Geometric Property
The following lemma serves as an important basis for the theoretical lower bounds in Theorems 3.2 and 3.5. The underlying intuition is that the orthosymmetry and convexity of the prior constraint $K$ imply the existence of vectors with small individual coordinates that nonetheless yield a significant aggregate norm. Such “spread-out” vectors, and the maximal norm they can attain, are effectively characterized by the Kolmogorov widths of $K$. Specifically, Lemma 3.1 guarantees the existence of a vector with a large $\ell_{2}$ norm but a small $\ell_{\infty}$ norm by leveraging the properties of QCO sets and the definition of the Kolmogorov widths.
@@
The proof of Lemma 3.1 is derived from the key properties of the QCO set $K$, combined with an application of the minimax theorem to a suitably truncated vector in $K^{2}$. The formal proof is provided in Appendix C.1. A similar argument also appears in Section 3.3.2 of Neykov, (2022).
Such existence in Lemma 3.1 is fundamental for the construction of the least favorable prior, where we define a uniform mixture over the coordinate-wise sign flips of this vector to induce a small total variation distance between the mixture and the null distribution $H_{0}$, thereby establishing one of the minimax lower bounds. This standard technique (see, e.g., LeCam, (1973)) can be adapted to our current problem with the prior constraint $K$ via Lemma 3.1.
Theorem 3.2 (First lower bound).
@@
Moreover, one possible explicit expressions of $c(\alpha)$ is
@@
For a fixed level $\alpha$, Theorem 3.2 establishes that if $\rho\lesssim\frac{(k_{1}^{\star})^{1/4}}{\sqrt{N}}\sigma$, no test can uniformly distinguish $H_{1}$ from $H_{0}$ with non-trivial power. As an immediate consequence, distinguishing $H_{0}$ and $H_{1}$ remains impossible in the presence of an adversary $\mathcal{C}$ under the same condition, since a valid strategy for $\mathcal{C}$ is simply applying the identity map (i.e., leaving the samples unperturbed).
Moreover, it is worth noting the marginal case when $k_{1}^{\star}=0$. Informally speaking, $k_{1}^{\star}=0$ means that the constraint set $K$ is insufficiently rich to support a signal that is statistically distinguishable from $H_{0}$. Since the definition of $k_{1}^{\star}$ only relies on $K$, NN, $\sigma$, and does not depend on the original samples $\widetilde{\mathbf{X}}$ or the observations $\mathbf{X}$, we shall assume in the following text that $D_{0}(K)=\sup\limits_{\theta\in K}\left\lVert\theta\right\rVert_{2}>\frac{1}{\sqrt{N}}\sigma$, and therefore $k_{1}^{\star}\geq 1$. Otherwise, the test will not be feasible in the first place. We kindly refer readers to Appendix C.2 for the proof of this argument and Theorem 3.2, which relies on the selected vector in Lemma 3.1 and Le Cam’s method of “fuzzy hypothesis”.
Recall that for uncontaminated observations without constraint, the critical value for $\left\lVert\mu\right\rVert_{2}$ is $\rho_{\text{critical}}\asymp\frac{d^{1/4}}{\sqrt{N}}\sigma$ (see Section 3.3.6 in Ingster and Suslina, (2003)). Theorem 3.2 generalizes this classic result in the sense that additional prior knowledge of the mean is allowed and the lower bound is refined as a consequence. Indeed, Theorem 3.2 recovers the classic result when we set $\mathcal{X}=\mathbb{R}^{d}$ (where $k_{1}^{\star}=d$ according to Example 2.5). Moreover, the theorem implies that when we have non-trivial constraint $K$ on $\mu$, one of the essential dimensional attributes of the problem is the finite $k_{1}^{\star}$ instead of $d$. (Recall that $k_{1}^{\star}<\infty$ is ensured even if $K\subset\ell_{2}$.)
@@ 3.1.2 Based on the Framework of Differential Privacy
Differential privacy was first introduced by Dwork et al., (2006) in the context of private data analysis. Formally, it requires the output distribution of a randomized algorithm to be insensitive to small perturbations in the input dataset. A key insight in this field is the strong connection between robust statistics and differential privacy; specifically, under mild conditions, robust estimators can be converted into differentially private algorithms using the propose-test-release framework. Subsequently, differential privacy was applied to the goodness-of-fit testing (also called identity testing) in statistics. Pioneering works in this direction include Gaboardi et al., (2016); Rogers and Kifer, (2017); Cai et al., (2017).
Definition 3.3 ($(\epsilon,\delta)$-differential privacy, Dwork et al., (2006)).
For a pair of non-negative real numbers $(\epsilon,\delta)$, a randomized algorithm $\mathcal{A}$ defined on the sample space $\Omega$ is a $(\epsilon,\delta)$-differentially private algorithm if for any $S\subset\text{range}(\mathcal{A})$ and any $X,X^{\prime}\in\Omega$ with the Hamming distance $d_{H}(X,X^{\prime})\leq 1$, we have
$\mathbb{P}\left(\mathcal{A}(X)\in S\right)\leq e^{\epsilon}\mathbb{P}\left(\mathcal{A}(X^{\prime})\in S\right)+\delta.$ (3.1)
When $\delta=0$, the notion is called pure differential privacy; when $\delta>0$, it is referred to as approximate differential privacy.
Our second lower bound relies on a fundamental coupling lemma: if two datasets $\mathbf{X}$ and $\mathbf{X}^{\prime}$, drawn from $\mathcal{U}$ and $\mathcal{V}$ respectively, can be coupled to overlap on a significant fraction of indices (i.e., $X_{i}=X^{\prime}_{i}$ for most $i$), then distinguishing between $\mathcal{U}$ and $\mathcal{V}$ is statistically hard.
Lemma 3.4 (Core lemma).
$\mathbb{E}_{(\mathbf{X},\mathbf{X}^{\prime})\sim(\mathcal{U},\mathcal{V})}d_{H}(\mathbf{X},\mathbf{X}^{\prime})\lesssim\epsilon N.$ (3.2)
The proof of Lemma 3.4 relies on a result established by Acharya et al., (2018), which states that if (3.2) holds for a $(0,\delta)$-differentially private algorithm, then necessarily $\epsilon N=\Omega(1/\delta)$. Furthermore, for any robust test attempting to distinguish between $\mathcal{U}$ and $\mathcal{V}$, one can construct a $(0,\delta)$-differentially private algorithm via the black-box robustness-to-privacy transformation proposed in Canonne et al., (2023) and Hopkins et al., (2023). Combining these two arguments yields a necessary condition for the expected Hamming distance between $\mathbf{X}$ and $\mathbf{X}^{\prime}$.
The following theorem, which establishes the second lower bound for the testing problem (1.2), builds upon Lemma 3.4 and the selected vector $\theta\in K$ associated with the second optimal dimension $k_{2}^{\star}$ (see Lemma 3.1). Specifically, let $\mathcal{U}=\mathbb{P}_{0}^{\otimes N}$ and $\mathcal{V}=\mathbb{P}_{1}^{\otimes N}$, where $\mathbb{P}_{0}$ denotes $\mathcal{N}(0,\sigma^{2})$ and $\mathbb{P}_{1}$ is the mixture of $\mathcal{N}(v,\sigma^{2})$ with $v$ drawn uniformly from the set
$V:=\left\{v\left\lvert\right.v_{i}=\gamma_{i}\theta_{i},\gamma_{i}\in\left\{-1,1\right\},i\geq 1,i\in I\right\}.$
We construct a chain of distributions interpolating between $\mathcal{U}$ and $\mathcal{V}$ based on the optimal coupling (Lemma B.5). Consequently, the computation of the expected Hamming distance is reduced to the sum of the total variation distances between adjacent distributions in the chain. We kindly refer readers to Appendix C.3 for the formal proofs of Lemma 3.4 and Theorem 3.5.
Theorem 3.5 (Second lower bound).
Consider the testing problem (1.2) under the strong $\epsilon$-contamination model with an adversary $\mathcal{C}$. If the signal strength satisfies $\rho\lesssim\frac{(k_{2}^{\star})^{1/4}\sqrt{\epsilon}\sigma}{N^{1/4}}$, then no test can simultaneously achieve type-I and type-II errors bounded by $0.1$. Formally, under this condition, we have
@@
where $A_{s}(0.1)$ denotes the set of tests with type-I error at most $0.1$, as defined in (1.3).
In Lemma 3.4 and Theorem 3.5, we fix the constant at $0.1$ for simplicity, though it can be replaced by any sufficiently small universal constant. In the high-dimensional setting where $d$ (and possibly $k_{2}^{\star}$) may substantially exceed $N$, one must ensure that
$\frac{(k_{2}^{\star})^{1/4}\sqrt{\epsilon}\sigma}{N^{1/4}}\leq D_{0}(K)=\sup\limits_{\theta\in K}\left\lVert\theta\right\rVert_{2}.$
If this condition is violated, the testing problem becomes statistically impossible.
@@ 3.1.3 Based on the Estimation Lower Bounds
The third lower bound, as established in Theorem 3.6 below, contrasts with Theorems 3.2 and 3.5 in that it is determined solely by the diameter of $K$ (denoted as $D_{0}(K)$), the fraction of contamination $\epsilon$, and the noise scale $\sigma$, while remaining independent of $N$.
Theorem 3.6 (Third lower bound).
For the signal detection problem (1.2) under the strong $\epsilon$-contamination model with an adversary $\mathcal{C}$, if $\epsilon\gtrsim\frac{1}{\sqrt{N}}$ and $\rho\lesssim\epsilon\sigma$, we always have
$\inf\limits_{\phi\in A_{s}}\sup\limits_{\mu\in K,\left\lVert\mu\right\rVert_{2}\geq\rho}\sup\limits_{\mathcal{C}}\mathbb{P}_{\mu}(\phi(\mathcal{C}(\widetilde{\mathbf{X}}))=0)=\varTheta(1),$ (3.3)
@@
(1). This lower bound is notable as it coincides with the minimax rate for the estimation problem under the strong $\epsilon$-contamination model (see, e.g., Prasadan and Neykov, (2025)). This finding stands in contrast to the conventional wisdom that testing is typically statistically easier than estimation (see Ingster, 1993a ; Baraud, (2002)). However, our subsequent analysis confirms that in this setting, the minimax rates for estimation and testing indeed match (or are of the same order) when certain conditions are met.
(2). The condition $\epsilon\gtrsim\frac{1}{\sqrt{N}}$ can be safely removed. See our discussions in Section 3.3.
Recall that $D_{0}(K)=\sup_{\theta\in K}\lVert\theta\rVert_{2}$ represents the radius of $K$ (i.e., the maximal distance from the origin). Theorem 3.6 implies that a necessary condition for the existence of a test with type-I and type-II errors uniformly bounded by $\alpha$ is that either $\rho\gtrsim\epsilon\sigma$ or $\epsilon=\mathcal{O}(1/\sqrt{N})$. Noting that the signal strength is inherently bounded by $\lVert\mu\rVert_{2}\leq D_{0}(K)$, it follows that if $\epsilon\gtrsim 1/\sqrt{N}$ and $\epsilon\sigma\gtrsim D_{0}(K)$, consistent testing is statistically impossible.
The proof of Theorem 3.6 is inspired by a similar lower bound established for the estimation problem under star-shaped constraints in Prasadan and Neykov, (2025). It relies on the construction of a specific distribution that admits two distinct mixture representations, with components corresponding to $H_{0}$ and $H_{1}$ respectively. Crucially, the adversary $\mathcal{C}$ can exploit this structural ambiguity to design a contamination strategy that renders the resulting distributions statistically indistinguishable when the lower bound condition is violated. The full proof is detailed in Appendix C.4.
@@ 3.2 Upper Bounds
The upper bounds are established constructively. Specifically, we design explicit testing procedures and identify the required signal strength $\rho=\lVert\mu\rVert_{2}$ that guarantees test validity (i.e., uniformly bounded type-I and type-II errors).
@@ 3.2.1 Theoretical Algorithm
Since the adversary $\mathcal{C}$ is allowed to modify only a limited number of samples, a natural strategy to mitigate the contamination is to identify subsets of $\mathbf{X}$ that exhibit a consistency property.
@@
[…]
The following theorem formally establishes the existence of a valid testing procedure under the condition (3.4), and consequently provides an upper bound for the testing problem (1.2). Specifically, we demonstrate that, when (3.4) holds, there exists a desired test $\phi_{e}$ and at least one consistent subset $\mathbf{X}_{S}$ with respect to $\phi_{e}$, upon which the entire testing task becomes feasible. To achieve this, we explicitly define an event $E_{e}$ of $\chi^{2}$-type, on which $\phi_{e}$ is determined, and prove that the original samples $\widetilde{\mathbf{X}}$ is consistent with high probability. The detailed proof is provided in Appendix C.5.
Theorem 3.8 (Theoretical upper bound).
Given a dataset $\mathbf{X}$ from the strong $\epsilon$-contamination model and a fixed error tolerance $\alpha$, for any $k\in[1,d]$ [^2] When $K\subset\ell_{2}$, we allow $1\leq k<\infty$., there exists a test $\phi_{e}$ for (1.2) such that the type-I and type-II errors are both uniformly less than $\alpha$ if $\rho^{2}\gtrsim E_{\text{raw}}^{2}(\epsilon,K,N,\sigma^{2},k)$ where $E_{\text{raw}}^{2}(\epsilon,K,N,\sigma^{2},k)$ is defined as
@@
We can further leverage the results of Theorem 3.8 by optimizing over $k$ and applying the definition of $k_{1}^{\star}$ and $k_{2}^{\star}$.
@@
Corollary 3.9 effectively selects the smaller of the optimal dimensions $k_{1}^{\star}$ and $k_{2}^{\star}$, and consequently determines the optimal projection on which $\phi_{e}$ is based. We show in Section 3.3 that the upper bound (3.5) matches the theoretical lower bounds up to a logarithmic factor of $\ln(1/\epsilon)$.
A significant improvement implied by Theorem 3.8 is that it does not require the asymptotic condition $\epsilon\rightarrow 0$, which is an essential assumption in many previous works. However, the test $\phi_{e}$ is limited by its reliance on an exhaustive search over all subsets of $\mathbf{X}$, resulting in exponential computational complexity with respect to $N$. This prohibitive cost renders the test computationally infeasible for large sample sizes. In the next section, motivated by the filtering method in Canonne et al., (2023), we propose a polynomial-time algorithm designed to handle the prior constraint $K$ in the presence of the adversary $\mathcal{C}$.
3.2.2 Polynomial-Time Algorithm
In this section, we develop a polynomial-time algorithm [^3] By polynomial-time”, we assume that computing the optimal projections (determining the optimal dimensions $k_{1}^{\star}$ and $k_{2}^{\star}$) can be performed in polynomial time, as is the case for any ellipsoid or hyperrectangle. to detect whether $\mu=\mathbf{0}$ under a slightly stronger condition than the bound (3.4). In contrast to the theoretical algorithm which requires an exhaustive search, the proposed algorithm leverages weights to adaptively filter potentially corrupted samples. Specifically, this filtering strategy and the algorithm are built and developed upon the framework of $\omega$-regularity, established in prior works (e.g., Canonne et al., (2023); Dong et al., (2019)).
Definition 3.10 ($\omega$-regularity).
(i), $\left|\sum\limits_{i\in S}\left\lVert Y_{i}\right\rVert_{2}^{2}-\left|S\right|k^{\star}\right|\leq c\beta_{1}$, (ii), $\left|\left\lVert\sum\limits_{i\in S}\sqrt{\omega}_{i}Y_{i}\right\rVert_{2}^{2}-\left\lVert\omega_{S}\right\rVert_{1}k^{\star}\right|\leq c\beta_{2}$, and (iii), $\left|\left\langle\sum\limits_{i\in S}\sqrt{\omega}_{i}Y_{i},\sum\limits_{j\in[N]}\sqrt{\omega}_{j}Y_{j}\right\rangle-\left\lVert\omega_{S}\right\rVert_{1}k^{\star}\right|\leq c\sqrt{N}\beta_{1}$,
where we recall that $\omega_{S}\in\mathbb{R}^{N}$ means the restriction of $\omega$ on the set $S$.
The intuition behind the polynomial-time algorithm parallels that of the theoretical construction, although establishing this connection relies on more advanced random matrix theory concerning (sub-)Gaussian random variables (see Appendix B for technical preliminaries). We first establish that the authentic samples satisfy the $\omega$-regularity condition for specified parameters $\beta_{1},\beta_{2}$ with high probability (setting the initial weights $\omega=\mathbf{1}_{[N]}$). Our analysis demonstrates that, under sufficient conditions on the signal strength $\rho=\lVert\mu\rVert_{2}$, the re-weighted samples effectively mimic the behavior of the unobserved authentic data $\widetilde{\mathbf{X}}$. Consequently, we are able to recover the test result corresponding to $\widetilde{\mathbf{X}}$. Such conditions are formalized in the following theorem.
Theorem 3.11 (Polynomial upper bound).